# Conceptualization, Data Curation

### v2

@@ 4.6 Transformers and Self-attention Networks
Transformer models were introduced by Vaswani et al. [87] in 2017 and are the foundation for many state-of-the-art NLP models. They consist of an encoder and a decoder, and the input and output are sequences. The innovation of this model is in the attention mechanism, which weights the importance of different samples in the input sequence, so the model can capture important information from the entire input sequence rather than a fixed window size. The transformer structure is shown in Figure 10.
Figure 10: Transformer structure [94]
@@
Fault happens when a variable deviates from the normal condition. Rapid changes can easily be captured by ML models, however, slow changes require a long time series, that can be analysed by attention-based models that have a global view of the whole process. Deep learning models like RNNs and CNNs have been widely applied for FDD of chemical processes, however, they are insufficient in extracting features of long-term dependencies. Self-attention mechanisms and transformer models can perform better for this goal. These models are primarily designed for natural language processing (NLP) tasks, however, time series data are very similar to text sequences [94]. One of the first researches that used transformers for chemical process fault detection is [121], they introduced a general transformer in which self-attention is replaced by graph attention, and it achieved an accuracy of 80.10% for 21 fault conditions. Wei et al. [94] proposed a Target Transformer algorithm, that has both self-attention and target-attention mechanisms. The model retains the original transformer’s encoder and applies target attention to the decoder to extract fault-related features. The average FDR for 20 faults of TEP is 90.39%. Zhou et al. [134] proposed an Industrial Process Optimization ViT (IPO-ViT) model based on Vision Transformer (ViT) for FDD. In this research, CNN which can extract local features and information is combined with multi-head attention which is famous for its global attention. FDR of the proposed model is increased compared to CNN and LSTM models, especially for faults 3,9 and 15. The average FDR for 20 faults of TEP is 94.08%.
@@
Zhang et al. [123], applied the self-attention (SA) mechanism to the gated recurrent unit (GRU), and introduced an encoder-decoder framework for FDD, where the encoder extracts the features and the decoder does the fault detection. The model is validated on TEP and the average recall rate for GRU-SA is 0.9504 which is better than GRU’s 0.9214. However, this approach can only be used when enough and reach fault samples are available. Xiong et al. [100], introduced a deep learning fault diagnosis method based on fully convolutional networks, long-short-term memory, and attention mechanism called attention-based LSTM and Fully Convolutional Network (ALSTM-FCN) for FDD. The CNN and LSTM layers are incorporated for their spatial and temporal feature extraction abilities respectively, and the attention mechanism is included for its ability to adaptively focus on the influence of important features. Dropout is used to prevent overfitting, and batch normalization to avoid the negative effects of covariate shift. Results show that the proposed method performs better in comparison to traditional CNN and LSTM models. The average accuracy for all 20 faults of TEP is 88.80% and for faults 3 and 9 less than 80%.
A summary of various algorithms that were successfully used for condition monitoring of TEP, along with their specifications is presented in Table LABEL:tab:models.
Table 3: Summary of Fault Detection Models, see 2 for fault types
@@
LSTM [118] Discrete wavelet transform and CNN for feature extraction and softmax classifier 1, 2, 4, 8, 10, 12, 13, 14, 17, and 18 98.73%
BiRNN [124] Bidirectional recurrent neural networks using gated recurrent unit (GRU) All 92.70%
LSTM + CNN [79] First LSTM and then CNN are used for feature extraction. The detection of faults 3, 9, and 15 are improved. All 92.06%
RNN [38] Using a 4 layer RNN, however, could not detect fault 15 All 92.50%
LSTM [38] Better performance compared to RNN speciallly for fault 15 All 95%
LSTM [122] Three layers LSTM are stacked All 99.49%
BiRNN [13] Bidirectional recurrent neural networks All 88.14%
GRU [58] The detetcion of fault 15 is enhanced compared to LSTM All 95.90%
CNN [96] CNN with dropout 20 faults 88.20%
CNN [120] FFT for feature extraction All faults except 3,9,15,16 96.90%
CNN [23] Handling multi time scale data using fixed multiple sampling method All 98.53%
CNN [30] Multiblock temporal convolutional network (MBTCN), dividing process variables into subblocks All faults except 3,9,15 96.20%
CNN [115] 1D CNN and Wavelet transform for feature extraction All 94.08%
CNN [90] Normalized CNN to improve the stability and speed of the training All 92.75%
CNN [28] Multiblock adaptive convolution kernel neural network (MBCKN) for large scale process using attention mechanism All 94.15%
CNN + LSTM [75] CNN is used to extract multi-level fault features followed by atcked LSTM solving the gradient disappearance problem. All 95.19%
CNN + LSTM [131] Extracts multiscale features for various operating conditions All faults except 3,9,15,21 98%
CNN + LSTM [15] Improved feature extraction and prevent model degradation All faults 98%
Autoencoder [37] Using wavelet activation function in the autoencoder All 95.86%
Autoencoder [112] Convolutional LSTM memory autoencoder (CLSTM-AE) and residual block for increasing the accuracy and make the training process more robust All 91%
Autoencoder [114] Manifold regularized stacked autoencoders (MRSAE) considering neighbourhood information needs a few labelled data All 91.42%
Autoencoder [125] Vector autoregressive model and 1-dimensional convolutional autoencoder for capturing the process dynamics All faults except 3, 9 and 15 96.80%
Autoencoder [62] Neuron-grouped stacked autoencoder (CG-SAE) as a hierarchical feature extractor for different groups of faults All faults except 3, 9 and 15 96.02%
Autoencoder [10] Orthogonal self-attentive variational autoencoder (OSAVA) to increase the interpretability All faults except 15, 16 94.70%
Autoencoder [40] Slow feature analysis-aided autoencoder (SFA-AE) to deal with slow variations in the data All faults except 15, 17 94.87%
Autoencoder [1] Combining LSTM and Supervised Autoencoder All 93.23%
Autoencoder [17] Variational recurrent autoencoder combines RNN and VAE All faults except 3,9,15 96.30%
Autoencoder [110] Supervised convolutional autoencoder (SCAE) for better feature extraction All faults except 3,9,15,16 96.19%
Autoencoder [89] Convolutional adversarial autoencoder (1DAAE) All 86%
Autoencoder [105] Nonlocal and global preserving stacked autoencoder (NLGPSAE) using a regularized objective function All 94.80%
@@
DBN [92] Extended deep belief network (EDBN) enhances the data quality during feature compression All faults except 3,9,15 94.31%
DBN [106] Adaptive lifting wavelet-DBN (ALW-DBN) to reduce the noise level All 97.33%
@@
DBN [104] Pairwise graph regularized deep belief network (PG-DBN) to monitor and control the feature manifold structure of the data All 93.87%
Transformer [121] Self attention transformer All 80.10%
@@
Transformer [94] Target transformer using both self-attention and target-attention mechanism 20 faults 90.39%
Transformer [134] Vision Transformer and CNN increasing the accuracy, especially for faults 3,9 and 15 20 faults 94.08%
@@
Attention mechanism [123] Combines self-attention (SA) mechanism to the gated recurrent unit (GRU) 20 faults 95.04%
@@
Attention mechanism [100] Attention-based LSTM and Fully Convolutional Network (ALSTM-FCN) with Dropout 21 faults 88.80%
@@
Recurrent Networks Process sequential or temporal data, modelling time dependencies Prone to gradient vanishing/exploding
@@
CNN Powerful feature extraction capability. Has denoising capabilities [49]. May need multiple layers and large labelled datasets to perform well
Autoencoder Feature extraction and dimension reduction ability. Unsupervised training. Robust against noise Pretraining needed
GAN Generate data to enhance the data imbalance problem. Can be used for semi-supervised classification [74] Stability problem in the training
@@
5 Discussion, Challenges and New Frontiers
This paper provides a summary of the most important deep learning algorithms that have been used for condition monitoring of industrial plants. Deep learning is a powerful tool for feature extraction and handling the large and complex data of chemical plants. LSTM and CNN models can find the hidden representations of the data. In general recurrent networks such as RNN, LSTM and GRU perform well on time series, sequential data and data with time dependencies. Transformers and attention mechanisms are also tools for analysing sequences by weighting the most important samples. Autoencoders reduce the dimension of the data and extract features. Different architectures and models, such as CNNs, can be used to design the encoder and decoder parts. Autoencoders are unsupervised learning methods and some of the variations like denoising Autoencoders are robust against noise. GANs are designed to generate data and can be used to enrich the dataset and increase the accuracy of the condition monitoring. A summary of the discussed methods and their advantages and limitations is provided in Table 4. The following part of this paper discusses challenges of data-driven condition monitoring. Some challenges like unlabeled, unseen or imbalanced data are related to the availability of the data, which can be handled by choosing the proper ML algorithm. Using AI and ML algorithms is consistently challenging due to their black-box nature. The system’s decisions are not easily interpretable for humans, posing difficulties in establishing trust and hindering real-world implementation. To address this problem, explainable AI, causal AI and ensemble learning approaches are introduced as a way to increase the accuracy and trust of the decision-making systems.
5.1 Unlabeled Data
@@
Collecting and labeling the data is challenging, especially in many industrial setups where even if the data is collected properly, labeling it may not be straightforward. Labelling the data is a time-consuming and expensive task that usually needs the supervision of an expert. Different approaches like unsupervised learning and Self-supervised learning have been used to deal with this problem. If only the normal data is labelled, anomaly detection methods can be useful.
@@
Most FDD methods fail when there is not enough labelled data. Peng et al. [69], proposed a method that uses a novel form of the bidirectional gated recurrent unit (BGRU), weighting each training example to overcome the class imbalance and cost-sensitive active learning for selecting the candidate instances to explore the unlabelled data. The proposed method is validated on TEP, by an imbalance ratio of 0.5 and the unlabelled proportion data of 0.6 and achieved the average FDR of 97.5%. Deep learning methods usually need large labelled datasets to achieve good performance. Self-supervised learning is an approach that can learn features from unlabeled data. Pöppelbaum et al. [72], proposed a novel self-supervised data augmentation technique based on SimCLR contrastive learning. The goal of contrastive learning is to compare different samples without any labels, the data itself provides supervision. In the SimCLR framework, multiple data augmentation methods for feature representation are used. By only considering 5% of the available training data, the classification accuracy is 80.80% for all TEP faults. To address situations where there are only a few labeled samples of normal data and the remaining available data, including abnormal samples, is unlabeled (positive-unlabeled (PU) learning), a common problem in the industry, a three-step high-fidelity PU (THPU) approach for FDD is introduced in [133]. The approach utilizes self-training nearest neighbors (STNN) to generate a positive dataset from unlabeled training data. Subsequently, a positive and negative data recognition (PNDR) algorithm, based on the SAE, augments the positive dataset and generates a negative dataset based on the unlabeled data. Finally, CNN is used as a binary classifier. With only one labelled normal sample, the FDR of the proposed method on TEP is 92.32% and will increase to 93.75% with 50 labelled normal samples. As Faults 3, 5, 9, 15 and 16 are very similar to the normal operational state, they were not considered in this study. A semi-supervised anomaly detection method composed of an autoencoder and LSTM model is introduced in [59], in this method, a control chart called Open Up is used for fault diagnosis. Open Up provides visual support that enables the interpretation and diagnosis of detected faults in a production system. The average accuracy of the model on a subset of TEP with 6.25% labelled data is 97.46%. Gorman et al. [25], proposed an interpretable anomaly detection method based on 1-Dimensional Convolutional Autoencoders (1dCAE) and Localised Reconstruction Error (LRE). In LRE each input channel is evaluated separately, and the sensors and data that lead to anomalies are identified, outperforming global reconstruction errors.
@@
Self-supervised contrastive learning extracts information from a large amount of unlabeled data. To enhance the performance of fault diagnosis in scenarios with numerous normal samples and only a few fault classes, a long-tailed distribution is employed in [65]. Progressively balanced supervised contrastive learning (PBS-SCL) is utilized to generate a balanced data batch, facilitating feature learning. Finally, a learnable linear classifier (LLC) is used for classification. The model was validated on a subset of TEP data. By having just 1,5,10,15, 20 samples of faults for each category and enough normal samples, the classification accuracy is 47.72%, 63.92%, 72.23%, 86.90%, and 86.59% respectively.
@@
Current zero-shot learning (ZSL) methods mostly focus on global level fault features; however, local level fault features that characterize the topological relationships between neighboring data points, are very important for generalization, to solve this problem a global–local attentionaware ZSL (GLA-ZSL) method is proposed in [85] consisting of a global prior refinement module using 2D-CNN for global features and local feature enhancement module based on the channel attention and 1D-CN, in addition a meta-learner is used for generalization. Human-defined fault attributes are used in this study.
5.2 Unseen Data

### v3

@@ 3.8 Transformers and Self-attention Networks
Transformer models were introduced by Vaswani et al. Vaswani et al. (2017) in 2017 and are the foundation for many state-of-the-art NLP models. They consist of an encoder and a decoder, and the input and output are sequences. The innovation of this model lies in its attention mechanism, which assigns varying weights to different samples in the input sequence. This enables the model to capture important information from the entire sequence, rather than just a fixed window size.
@@
Fault happens when a variable deviates from the normal condition. Rapid changes can easily be captured by ML models; however, slow changes require a long time series, which can be analysed by attention-based models that have a global view of the whole process. Deep learning models like RNNs and CNNs have been widely applied for FDD of chemical processes; however, they are insufficient in extracting features of long-term dependencies. Self-attention mechanisms and transformer models can perform better for this goal. These models are primarily designed for natural language processing (NLP) tasks; however, time series data are very similar to text sequences Wei et al. (2022). One of the first studies that used transformers for chemical process fault detection is Zhang et al. (2022), they introduced a general transformer in which self-attention is replaced by graph attention, and it achieved an accuracy of 80.10% for 21 fault conditions. Wei et al. Wei et al. (2022) proposed a Target Transformer algorithm that has both self-attention and target-attention mechanisms. The model retains the original transformer’s encoder and applies target attention to the decoder to extract fault-related features. The average FDR for 20 faults of TEP is 90.39%. Zhou et al. Zhou et al. (2023) proposed an Industrial Process Optimization ViT (IPO-ViT) model based on Vision Transformer (ViT) for FDD. In this research, a CNN, which can extract local features and information, is combined with multi-head attention, which is famous for its global attention. FDR of the proposed model is increased compared to CNN and LSTM models, especially for faults 3,9, and 15. The average FDR for 20 faults of TEP is 94.08%.
@@
Zhang et al. Zhang and Xiong (2019), applied the self-attention (SA) mechanism to the gated recurrent unit (GRU), and introduced an encoder-decoder framework for FDD, where the encoder extracts the features and the decoder does the fault detection. The model is validated on TEP, and the average recall rate for GRU-SA is 0.9504, which is better than GRU’s 0.9214. However, this approach can only be used when enough and rich fault samples are available. Xiong et al. Xiong et al. (2023) introduced a deep learning fault diagnosis method based on fully convolutional networks, long-short-term memory, and attention mechanism called ALSTM-FCN for FDD. The CNN and LSTM layers are incorporated for their spatial and temporal feature extraction abilities, respectively, and the attention mechanism is included for its ability to adaptively focus on the influence of important features. Dropout is used to prevent overfitting, and batch normalization is used to avoid the negative effects of covariate shift. Results show that the proposed method performs better in comparison to traditional CNN and LSTM models. The average accuracy for all 20 faults of TEP is 88.80%, and for faults 3 and 9, it is less than 80%.
Attention mechanisms and transformers are emerging tools for long-sequence modeling in industrial CM, outperforming traditional RNN/CNN hybrids when capturing slow drifts and long-range correlations. Vision Transformer–CNN hybrids (IPO-ViT) and target-attention transformers achieve improved detection for challenging faults (3, 9, 15) by combining local feature extraction with global attention. However, these models are data-hungry and can overfit when fault samples are scarce. In such cases, simpler attention-augmented RNNs or CNNs often provide a better accuracy–complexity trade-off. Transformer adoption in CM is still in its early stages, and future work will likely focus on reducing computational demands while improving robustness to noise. A summary of these methods is shown in 6.
Table 6: Summary of DBN and transformers and Attention networks. See Table 2 for fault types.
@@
DBN Wang et al. (2020) Extended deep belief network (EDBN) enhances the data quality during feature compression All faults except 3,9,15 94.31%
DBN Yao et al. (2021) Adaptive lifting wavelet-DBN (ALW-DBN) to reduce the noise level All 97.33%
@@
DBN Yang et al. (2021) Pairwise graph regularized deep belief network (PG-DBN) to monitor and control the feature manifold structure of the data All 93.87%
Transformer Zhang et al. (2022) Self attention transformer All 80.10%
@@
Transformer Wei et al. (2022) Target transformer using both self-attention and target-attention mechanism 20 faults 90.39%
Transformer Zhou et al. (2023) Vision Transformer and CNN increasing the accuracy, especially for faults 3,9, and 15 20 faults 94.08%
@@
Attention mechanism Zhang and Xiong (2019) Combines self-attention (SA) mechanism with the gated recurrent unit (GRU) 20 faults 95.04%
@@
Attention mechanism Xiong et al. (2023) Attention-based LSTM and Fully Convolutional Network (ALSTM-FCN) with Dropout 21 faults 88.80%
4 Discussion, Challenges and New Frontiers
This study provides a summary of the most important deep learning algorithms that have been used for condition monitoring of industrial plants. Deep learning is a powerful tool for feature extraction and handling the large and complex data of chemical plants. LSTM and CNN models can find the hidden representations of the data. In general, recurrent networks such as RNN, LSTM, and GRU perform well on time series, sequential data, and data with time dependencies. Transformers and attention mechanisms are also powerful tools for analyzing sequences by weighting the most important samples. Autoencoders reduce the dimension of the data and extract features. Different architectures and models, such as CNNs, can be used to design the encoder and decoder parts. Autoencoders are unsupervised learning methods, and some of the variations, like denoising Autoencoders, are robust against noise. GANs are designed to generate data and can be used to enrich the dataset and increase the accuracy of the condition monitoring. A summary of the discussed methods and their advantages and limitations is provided in Table 7. The following part of this paper discusses the challenges of data-driven condition monitoring. Some challenges, such as unlabeled, unseen, and imbalanced data, are related to the availability of the data, which can be handled by choosing the proper ML algorithm. Using AI and ML algorithms is consistently challenging due to their black-box nature. The system’s decisions are not easily interpretable for humans, posing difficulties in establishing trust and hindering real-world implementation. To address this problem, explainable AI, causal AI, and ensemble learning approaches are introduced as a way to increase the accuracy and trust of the decision-making systems.
@@ 4 Discussion, Challenges and New Frontiers
Recurrent Networks Process sequential or temporal data, modeling time dependencies Prone to gradient vanishing/exploding
@@
CNN Powerful feature extraction capability. Has denoising capabilities Lu et al. (2017). May need multiple layers and large labeled datasets to perform well
Autoencoder Feature extraction and dimension reduction ability. Unsupervised training. Robust against noise Pretraining is needed
GAN Generate data to enhance the data imbalance problem. Can be used for semi-supervised classification Qin and Zhao (2022b) Stability problem in the training
@@
4.1 Data Challenges
This section addresses common data challenges in industrial environments, namely unlabeled data, unseen fault scenarios, and class imbalance. These issues, illustrated in Figure 4, often limit the performance of conventional ML and DL algorithms. To address them, we review advanced strategies specifically designed to mitigate these challenges and thereby improve the robustness and reliability of condition monitoring in industrial applications.
Figure 4: Common data challenges in industrial condition monitoring
4.1.1 Unlabeled Data
@@
Collecting and labeling data can be quite challenging, especially in industrial settings. Even when data is properly collected, labeling it may not be straightforward. This labeling process is often time-consuming and expensive, typically requiring the supervision of an expert. Various approaches, such as unsupervised learning and self-supervised learning, have been employed to address this issue. If only the normal data is labeled, anomaly detection methods can be effective.
@@
Most FDD methods fail when there is not enough labeled data. Peng et al. Peng et al. (2020b) proposed a method that uses a novel form of the bidirectional gated recurrent unit (BGRU), weighting each training example to overcome the class imbalance and cost-sensitive active learning for selecting the candidate instances to explore the unlabeled data. The proposed method is validated on TEP, with an imbalance ratio of 0.5 and the unlabeled proportion data of 0.6, and achieved an average FDR of 97.5%. Deep learning methods usually need large labeled datasets to achieve good performance. Self-supervised learning is an approach that can learn features from unlabeled data. Pöppelbaum et al. Pöppelbaum et al. (2022) proposed a novel self-supervised data augmentation technique based on SimCLR contrastive learning. The goal of contrastive learning is to compare different samples without any labels; the data itself provides supervision. In the SimCLR framework, multiple data augmentation methods for feature representation are used. By only considering 5% of the available training data, the classification accuracy is 80.80% for all TEP faults. To address situations where there are only a few labeled samples of normal data and the remaining available data, including abnormal samples, is unlabeled (positive-unlabeled (PU) learning), a common problem in the industry, a three-step high-fidelity PU (THPU) approach for FDD is introduced in Zheng and Zhao (2022). The approach utilizes self-training nearest neighbors (STNN) to generate a positive dataset from unlabeled training data. Subsequently, a positive and negative data recognition (PNDR) algorithm, based on the SAE, augments the positive dataset and generates a negative dataset based on the unlabeled data. Finally, CNN is used as a binary classifier. With only one labeled normal sample, the FDR of the proposed method on TEP is 92.32% and will increase to 93.75% with 50 labeled normal samples. Faults 3, 5, 9, 15, and 16, which are very similar to the normal operational state, were not considered in this study. A semi-supervised anomaly detection method composed of an autoencoder and an LSTM model is introduced in Morales-Forero and Bassetto (2019). In this method, a control chart called Open Up is used for fault diagnosis. Open Up provides visual support that enables the interpretation and diagnosis of detected faults in a production system. The average accuracy of the model on a subset of TEP with 6.25% labeled data is 97.46%. Gorman et al. Gorman et al. (2022) proposed an interpretable anomaly detection method based on 1-Dimensional Convolutional Autoencoders (1dCAE) and Localised Reconstruction Error (LRE). In LRE, each input channel is evaluated separately, and the sensors and data that lead to anomalies are identified, outperforming global reconstruction errors. Chadha et al. Chadha et al. (2021) used a convolutional autoencoder for handling unlabeled data. The latent features are split into discriminative and reconstructive, and an auxiliary loss based on k-means clustering is used for the discriminatory latent variables. The classification average accuracy for 3, 9, 15, and 21 faults, which are very hard to detect, is 50.04%. Usually, there are not enough labeled samples, so semi-supervised learning methods are practical as they use unlabeled data to help with the limited labeled data during model training. A consistency regularization autoencoder (CRAE) framework is introduced in Ma et al. (2022), which captures correlations from labeled and unlabeled data. Previous samples are augmented with the current samples to increase the fault detection speed and accuracy. With just 20% of the dataset’s labels, the average fault detection accuracy for all faults in the TE dataset, excluding faults 3, 9, and 15, is 92.95%. To utilize unlabeled data more efficiently, Zhang et al. Zhang and Qiu (2022b), combined LSTM for temporal feature extraction and ladder autoencoder (LAE) for semi-supervised learning. LSTM-LAE achieves the interpretability to extract fault-relevant process variables. By just having the labels of 11% of the dataset, this model achieved the average accuracy of 95.9% for all faults except 3,13, 15, and 16.
@@
Self-supervised contrastive learning extracts information from a large amount of unlabeled data. To enhance the performance of fault diagnosis in scenarios with numerous normal samples and only a few fault classes, a long-tailed distribution is employed in Peng et al. (2022a). Progressively balanced supervised contrastive learning (PBS-SCL) is utilized to generate a balanced data batch, facilitating feature learning. Finally, a learnable linear classifier (LLC) is used for classification. The model was validated on a subset of TEP data. With 1.5, 10, 15, 20 samples of faults for each category and sufficient normal samples, the classification accuracy is 47.72%, 63.92%, 72.23%, 86.90%, and 86.59%, respectively.
@@
Current zero-shot learning (ZSL) methods mostly focus on global level fault features; however, local level fault features that characterize the topological relationships between neighboring data points, are very important for generalization, to solve this problem a global–local attentionaware ZSL (GLA-ZSL) method is proposed in Tang et al. (2025) consisting of a global prior refinement module using 2D-CNN for global features and local feature enhancement module based on the channel attention and 1D-CN, in addition a meta-learner is used for generalization. Human-defined fault attributes are used in this study. A dual-network autoencoder-based adversarial domain adaptation with Wasserstein divergence was introduced in Yang et al. (2024) to diagnose unlabeled data. A dual-network autoencoder combining a convolutional neural network (CNN) and long short-term memory (LSTM) is used to extract both local deep features and temporal information. The autoencoder performs unsupervised reconstruction of source domain data, ensuring high classifier accuracy. An adversarial training process is employed with a domain discriminator to guide the feature extractor in learning domain-invariant features, minimizing the Wasserstein distance to reduce the difference in feature distribution across domains. Additionally, Wasserstein divergence is introduced to the adversarial process to enhance the stability of the training.
4.1.2 Unseen Data